OpenStudy (maheshmeghwal9):

What is the value of i^i? Where i=iota, a complex no. (0,1). Please tell!

OpenStudy (maheshmeghwal9):

the answer is \[e^(-pi/2)\]

OpenStudy (maheshmeghwal9):

e raised to power - pi/2

OpenStudy (maheshmeghwal9):

i have just done upto \[e^\ln(i^i) \]

OpenStudy (maheshmeghwal9):

e raised to the power of log of i squared to the base e

OpenStudy (maheshmeghwal9):

now what should i do from here?

OpenStudy (anonymous):

\[i = \cos pi/2 + i (\sin pi/2)\] Therefore \[\large i = e^{i(pi/2)}\] Now\[\huge i^i = e^{(i*i(pi/2))} = e ^ {-pi/2}\]

OpenStudy (anonymous):